Steam at 30 bar is contained in a cylinder behind a piston the initial volume being 0.08067 m'kg calculate the heat flow to or form cylinder and the work done by the steam a) According to the low pys - C to final volume 0.2 m / kg. b) According to a linear law to final dry saturated volume of 0.1047m / kg



1
Expert's answer
2021-05-06T07:41:55-0400

"A=p\Delta V,"

a)

"A_1=p(V_2-V_1)=358~kJ,"

b)

"A_2=p(V_3-V_1)=72~kJ."